Pam is 14 and her dad is 37. In how many years will pans dad be twice as old as she will be???
Answer by jim_thompson5910(35256)

Let x = number of years


you're adding x to each age to get


Pams age after x years: 14+x
Dad's age after x years: 37+x

Since the question is "In how many years will pans dad be twice as old as she will be?", this means that the equation is then


2(14+x)=37+x


Now solve for x


2(14+x)=37+x


28+2x=37+x


2x=37+x-28


2x-x=37-28


x=37-28


x=9


So the solution is x=9


This means that it will take 9 years.